OpenStudy (highschoolmom2010):

A sphere has center (0, 0, 0) and radius 5. Tell whether each of the following points is inside, outside, or on the sphere. a. A(0, -3, 4) b. B(1, -1,-1) c. C(4, -6, -10)

OpenStudy (highschoolmom2010):

which equation would i need to use

OpenStudy (highschoolmom2010):

what are x,y,z

OpenStudy (anonymous):

For this problem you need to calculate the distance from the origin for each point. If it is less or more than 5

OpenStudy (anonymous):

Do you know how to do that for a 3 dimensional space?

OpenStudy (anonymous):

His method was over-complicated.

OpenStudy (highschoolmom2010):

um not a clue how to do that

OpenStudy (anonymous):

So here is the general formula for the distance between 2 point P(p1,p2,p3) and Q(q1,q2,q3) in 3D:

OpenStudy (anonymous):

\[d(p,q)=\sqrt{(p _{1}-q _{1})^2+(p _{2}-q _{2})2+(p _{3}-q _{3})^2}\]

OpenStudy (anonymous):

Here luckily P is (0,0,0)

OpenStudy (anonymous):

Q will be the 3 points given. When you put these in you will get a number, If it is =5 it is on the sphere, less than 5 inside, more than 5 outside. Hope this helped

OpenStudy (anonymous):

I will do B for you as that is the easiest: distance d(0,B)=\[\sqrt{(0-1)^2+(0-(-1))^2+(0-(-1))^2}=\sqrt{3}\]

OpenStudy (anonymous):

So it is inside the sphere

OpenStudy (anonymous):

A will be on the sphere, C will be outside. But you will have to do the calculation :)

OpenStudy (highschoolmom2010):

yes thanks for explaining